Using Maxwell’s equations in free space, derive the wave equation for the z-component of
the electric field vector.
1
Expert's answer
2019-03-27T04:39:51-0400

Let us consider the Maxwell's equations in free space (in Gauss units)


"\\nabla \\cdot \\vec{E} = \\rho\\\\\n\\nabla \\cdot \\vec{H} = 0\\\\\n\\nabla \\times \\vec{E} = - \\frac{1}{c} \\frac{\\partial \\vec{H}}{\\partial t}\\\\\n\\nabla \\times \\vec{H} = \\frac{1}{c} \\frac{\\partial \\vec{E}}{\\partial t}\t+ \\frac{4 \\pi}{c} \\vec{j}"

After calculating the curl from the left side of the third equation, we obtain:


"\nabla \times (\nabla \times \vec{E}) = \nabla(\nabla \cdot \vec{E}) - \Delta \vec{E} = \nabla \rho - \Delta \vec{E}"

Performing this procedure over the right side of the same equation and taking into account that partial derivative over time commutes with nable operator, we obtain:


"\\nabla \\times \\left( - \\frac{1}{c} \\frac{\\partial \\vec{H}}{\\partial t} \\right) = - \\frac{1}{c} \\frac{\\partial}{\\partial t} \\left( \\frac{1}{c} \\frac{\\partial \\vec{E}}{\\partial t}\t\n \t+ \\frac{4 \\pi}{c} \\vec{j} \\right) = - \\frac{1}{c^2} \\frac{\\partial^2 \\vec{E}}{\\partial t^2} - \\frac{4 \\pi}{c^2} \\frac{\\partial \\vec{j}}{\\partial t}"

Finally, putting the expressions together, we deduce:


"\Delta \vec{E} - \frac{1}{c^2} \frac{\partial^2 \vec{E}}{\partial t^2} = \nabla \rho + \frac{4 \pi}{c^2} \frac{\partial \vec{j}}{\partial t}"

In order to get the equation for z-component, one should project this expression onto the z-axis:


"\Delta E_z - \frac{1}{c^2} \frac{\partial^2 E_z}{\partial t^2} = \frac{\partial \rho}{\partial z} + \frac{4 \pi}{c^2} \frac{\partial j_z}{\partial t}"